### v2

# Training-free Camera Control for Video Generation

Abstract
We propose a training-free and robust solution to offer camera movement control for off-the-shelf video diffusion models. Unlike previous work, our method does not require any supervised finetuning on camera-annotated datasets or self-supervised training via data augmentation. Instead, it can be plugged and played with most pretrained video diffusion models and generate camera controllable videos with a single image or text prompt as input. The inspiration of our work comes from the layout prior that intermediate latents hold towards generated results, thus rearranging noisy pixels in them will make output content reallocated as well. As camera move could also be seen as a kind of pixel rearrangement caused by perspective change, videos could be reorganized following specific camera motion if their noisy latents change accordingly. Established on this, we propose our method CamTrol, which enables robust camera control for video diffusion models. It is achieved by a two-stage process. First, we model image layout rearrangement through explicit camera movement in 3D point cloud space. Second, we generate videos with camera motion using layout prior of noisy latents formed by a series of rearranged images. Extensive experiments have demonstrated the robustness our method holds in controlling camera motion of generated videos. Furthermore, we show that our method can produce impressive results in generating 3D rotation videos with dynamic content. Project page at https://lifedecoder.github.io/CamTrol/.
1 Introduction
Figure 1: Training-free video camera control by CamTrol. CamTrol could handle basic camera motions, hybrid motions and complicated trajectories with precise camera coordinates. Besides, CamTrol produces impressive results on generating 3D rotation videos in various styles. Note that all human faces presented in this paper are synthesized using text-to-image models.
As a more appealing and content-richer modality, videos differ from images by including an extra temporal dimension. This temporal aspect provides increased versatility for depicting diverse and dynamic movements, which can be decomposed into object motion, background transitions and perspective changes. Recent years have witnessed the rapid development and splendid breakthrough of video generation with text prompt or images as input instructions li2023videogen ; hong2022cogvideo ; ho2022imagenvideo ; luo2023videofusion ; zeng2023PixelDance ; SVD ; sora ; ge2023preserve ; fei2023DysenVDM , and demonstrated the inestimable potential of diffusion models to synthesis realistic videos. While these video generation models have made progress in generating videos with highly dynamic objects and backgrounds zeng2023PixelDance ; SVD ; li2023videogen , most of them fail to provide camera control for the generated videos.
The difficulty of controlling camera trajectory in videos primarily arises from two aspects. The initial challenge lies in the inadequacy of annotated data. Most video annotations lack of descriptions, especially precise descriptions of video’s camera movements. As a result, video generation models trained on these data often fail to interpret text prompts related to camera motions and generate correct outputs. One solution to mitigate the data insufficiency problem is to mimic videos with camera movements through simple data augmentation yang2024direct . However, these methods could only handle simple camera motions like zoom or truck, and have trouble in dealing with more complicated ones. The second challenge lies in the effort of additional finetuning required for controlling camera movements. As camera trajectories could be sophisticated, they sometimes cannot be accurately elaborated using naïve text prompts alone. Common solutions wang2023motionctrl ; he2024cameractrl proposed to embed camera parameters into diffusion models through learnable encoders and perform extensive finetuning on large-scale datasets with detailed camera trajectories. However, such datasets like RealEstate10k realestate10k and MVImageNet yu2023mvimgnet are intensively limited in scale and diversity due to the difficulty associated with data collection, in this way, these finetuning methods demand substantial resources but exhibit limited generalizability to other types of data. Lack of annotations and heavy finetuning effort make camera control a challenging task in video generations.
In this work, we attempt to address these issues through a training-free solution to offer camera control for off-the-shelf video diffusion models. We begin by introducing two core observations underpin that video diffusion models can achieve camera movement control in a training-free manner. First, we find that base video models could produce results with rough camera moves by integrating specific camera-related text into input prompts, such as camera zooms in or camera pans right. This simple implementation, though not very accurate and always leads to static or wrong motions, shows the natural prior knowledge learnt by pretrained models about following different camera trajectories. The other observation is the effectiveness video models have exhibited in adapting to 3D generation tasks. Recent works sv3d ; im3d ; shi2023mvdream find that leveraging pretrained video models as initialization helps drastically improve the performance of multi-view generations, demonstrating their strong ability of handling perspective change. The two crucial observations reveals the hidden power of video models for camera motion control, thus, we seek to find a way to evoke this innate ability, as it already exists in the model itself.
We propose CamTrol, which offers camera control for off-the-shelf video diffusion models in a training-free but robust manner. CamTrol is inspired by the layout prior that noisy latents hold towards generation results: As pixels in noisy latents change their positions, corresponding rearrangement will also occur to the output and leads to layout modification. Considering camera moves could also be seen as a kind of layout rearrangement, this prior can serve as an efficacious hint providing video model with information of specific camera motions. Specifically, CamTrol consists of a two-stage procedure. In stage I, explicit camera movements are modeled in 3D point cloud representations and produce a series of rendered images indicating specific camera movements. In stage II, layout prior of noisy latents are utilized to guide video generations with camera movements. We conduct extensive experiments to validate the effectiveness of our proposed CamTrol. Both quantitative and qualitative results demonstrate the robustness of CamTrol as a useful tool of controlling camera motion for video diffusion models. Furthermore, we show that CamTrol  produces impressive results of dynamic 3D rotation videos in various styles.
2 Related Work
Camera Control for Video Generation
While methods aim for controlling video foundation models constantly emerge ma2024magic ; liu2023video ; feng2023ccedit , there are few works explore how to manipulate camera motion of generated videos. Earlier work hao2018controllable controls motion trajectory via warping image through densified sparse flow and pixel fusion, similar ideas also appear later in chen2023motion ; yin2023dragnuwa . Besides utilizing optical flow, two main techniques for implementing video camera control are via self-supervised augmentation or additional finetuning. yang2024direct disentangles object motion with camera movement and incorporates extra layers to embed camera motions, where model is trained in a self-supervised manner by augmenting input videos to stimulate simple camera movements. he2024cameractrl and wang2023motionctrl train an additional camera encoder and integrate the output into temporal attention layers of U-Net. guo2023animatediff learns new motion pattern via LoRA hu2021lora and finetuning with multiple reference videos.
Noise Prior of Latents in Diffusion Model
One of the most natural advantages of diffusion model comes from its pixel-wise noisy latents formed during denoising process. These latents hold strong causuality towards output and directly determine what the result looks like, meanwhile have robust error-resilience as they are perturbed by Gaussian noises across different scales. Numerous work have exploited the convenience of this noise prior to attain controllable generation, such as image-to-image translation meng2021sdedit , pixel-level manipulation nichol2022glide ; avrahami2022blended , image inpainting lugmayr2022repaint and semantic editing choi2021ilvr ; kim2022diffusionclip ; hou2024high . Recent research has shown that even sampled from Gaussian distribution, the initial noise of diffusion process still have significantly influence to the layout of generated contents mao2023guided . In other work, noise prior is used to guarantee temporal consistency among video frames luo2023videofusion , or to trade-off between fidelity and diversity of image editing kim2022diffusionclip .
Video Model for 3D Generation
Similar to how most video generation models using the groundwork laid by image foundation models blattmann2023alignVideo-LDM ; esser2023GEN1 ; singer2022make ; tuneavideo , training of 3D generation model also relies heavily on pretrained 2D video models sv3d ; im3d ; shi2023mvdream ; chen2024v3d ; han2024vfusion3d . These methods either finetunes with rendered videos directly SVD ; chen2024v3d ; im3d ; han2024vfusion3d , or adds camera embedding for each view as extra condition sv3d ; shi2023mvdream . Video foundation models have shown to be particularly beneficial in generating consistent multi-view rendering of 3D objects, demonstrates their inherent abundant prior knowledge for handling camera pose change.
3 Training-free Camera Control for Video Generation
CamTrol takes two stages to evoke the innate camera control ability hidden in base video models. In Sec. 3.1, we will describe how to model explicit camera motion in point cloud space. In Sec. 3.2, we will elaborate on the camera-controllable video generation with the guidance of noise layout prior.
Figure 2: Pipeline of CamTrol. In stage I, camera movements are modeled through explicit 3D point cloud . In stage II, layout prior of noisy latents are utilized to guide video generation.
3.1 Camera Motion Modeling
To evoke pretrained video diffusion model’s ability of dealing with camera perspective changes, hints of camera motion should be injected to diffusion model in a proper way. While simply concatenating camera trajectories with text prompt is incomprehensible for original model, previous works wang2023motionctrl ; he2024cameractrl introduce additional embedder to encode camera parameters and finetune with limited annotated data realestate10k ; yu2023mvimgnet , which are data-hungry yet lack of generalization ability. Other methods yang2024direct construct camera motions by self-supervised augmentations, but could only handle a few easy camera controls. Thus, we seek a more efficient and robust way to guide the model towards camera controllable.
Considering perspective change of video is originally caused by camera movements in 3D space, we resort to 3D representation for providing explicit motion hints to video diffusion models. Specifically, we choose point cloud as the intermediate representation, in which space we can expediently manipulate camera poses and positions for simulating diverse camera movements. Besides explicit camera modeling, introducing point cloud brings extra benefits: The first is its data-efficiency. By utilizing inpainting techniques, only one single input image is required for the whole point cloud reconstruction, this sidesteps the effort of large-scale finetuning. Second, consistency between multi-view renderings can be easily ensured, as the known points will remain unchanged once the reconstruction finished.
Point Cloud Initialization
We start by lifting the pixels in input image plane to 3D point cloud representations. In practice, the input image can be either user-defined or created by image generators like Stable Diffusion rombach2022high . Given an input image $\mathbf{I}_{0}\in\mathbb{R}^{3\times H\times W}$, we first estimate its depth map $\mathbf{D}_{0}$ using off-the-shelf monocular depth estimator ZeoDepth bhat2023zoedepth . By combining image and its depth map, point cloud $\mathcal{P}_{0}$ can be initialized as:
$\mathcal{P}_{0}=\phi([\mathbf{I}_{0},\mathbf{D}_{0}],\mathbf{K},\mathbf{P}_{0}),$ (1)
where $\phi$ denotes the mapping function from RGBD to 3D point cloud, $\mathbf{K}$ and $\mathbf{P}_{0}$ represent camera’s intrinsic and extrinsic matrices set by convention chung2023luciddreamer as they’re usually intractable.
Camera Trajectories
To get consistent images from multiple viewpoints, we set camera motion as a pre-defined trajectory of extrinsic matrix $\left\{\mathbf{P}_{1},...,\mathbf{P}_{N-1}\right\}$, each of which includes a rotation matrix and translation matrix representing camera’s pose and position. At each step $i$, we project the point cloud back to camera plane using function $\psi$ and get a rendered image with perspective change: $\mathbf{I}_{i}=\psi(\mathcal{P}_{i},\mathbf{K},\mathbf{P}_{i})$. By calculating extrinsic matrices of corresponding movement, we obtain a series of camera motions including zoom, tilt, pan, pedestal, truck, roll and rotate, enabling flexible camera movements. Detailed definitions of these movements are elaborated in Appendix A.1. By combining basic trajectories, hybrid camera movements can be attained and produce videos with cinematic charm. What’s more, benefit from explicit camera motion modeling, our method could support trajectories with precise coordinates, which means it can generate videos with any complicated camera motion(Fig. 1).
Multi-view Rendering
When perspective changes, there can be vacancies appear as some areas are unoccupied within the point cloud. To get more reasonable results, we employ image inpainting model rombach2022high to fill up the holes for new renderings, with a mask distinguishing the known points from nonexistent ones. After inpainted in 2D space, image is lifted again onto 3D space and gradually complete the whole point cloud representation. During this process, misalignment between adjacent views may occur since depth estimator only estimates relative depth, further leads to inconsistency of rendered images. We adopt depth coefficient optimization chung2023luciddreamer to avoid this situation, which can be formed as:
$d_{i}=\operatorname*{argmin}_{d}\left(\sum_{M}\left\|\phi([\tilde{\mathbf{I}}_{i},d\tilde{\mathbf{D}}_{i}],\mathbf{K},\mathbf{P}_{i})-\mathcal{P}_{i-1}\right\|\right),$ (2)
where $\tilde{\mathbf{I}}_{i}$ and $\tilde{\mathbf{D}}_{i}$ refer to the inpainted image and its depth map respectively, $d_{i}$ denotes depth coefficient to be optimized, and $M$ refers to the overlapping region between $\mathcal{P}_{i}$ and $\mathcal{P}_{i-1}$, as other areas are not shared for calculating $\ell_{1}$ loss.
Thus, we get a set of images refer to the input and indicate specific camera movement:
$\left\{\mathbf{I}_{0},...,\mathbf{I}_{N-1}\right\}=\left\{\psi(\mathcal{P}_{i},\mathbf{K},\mathbf{P}_{i})|i\in\left[0,N-1\right]\right\}.$ (3)
3.2 Layout Prior of Noise
With camera motion modeling, we obtain a sequence $\mathbf{V}_{0}=\left[\mathbf{I}_{0},...,\mathbf{I}_{N-1}\right]\in\mathbb{R}^{N\times 3\times H\times W}$ of rendered images adhering to a specific camera trajectory. Note that quality of rendered images are not perfect as single input image only leads to sparse point cloud reconstruction, besides, these renderings are static, thus they could not use directly as video frames. To form an ideal video, we need to find a way that satisfies the following three requirements: 1) camera motions should be maintained; 2) video should be encouraged with more dynamics; and 3) quality imperfection should be compensated.
Camera Motion Inversion
Recent work on diffusion models have demonstrated the strong controllability of its noisy latents meng2021sdedit ; mao2023guided , the causality and error-resilience they hold towards final output make them a convenient yet powerful tool for controllable generation of diffusion models. Particularly for initial noise, even sampled from Gaussian distribution, it still have significant influence on the layout of generated image mao2023guided , so that rearranging the noise pixels will make content in output relocate as well. For instance, if all pixels in initial noise shift to right by a certain distance, it is likely that generated output reflect a similar shift. This reminds us that the impact of camera movement on images could also be regarded as a kind of layout rearrangement, where pixels change their positions caused by viewpoint change. In a similar way, videos can be reorganized following camera motion if their noisy latents change accordingly.
Inspired by this, we first construct a series of noisy latents indicating corresponding camera movements. It can be intuitively done by employing diffusion model’s inversion process on rendered image sequence $\mathbf{V}_{0}$. Latent at timestep $t_{0}$ can be calculated as:
$\mathbf{V}_{t_{0}}=\sqrt{\bar{\alpha}_{t_{0}}}\mathbf{V}_{0}+\sqrt{1-\bar{\alpha}_{t_{0}}}\epsilon,\quad\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I}),$ (4)
where $\bar{\alpha}_{t}$ are variances used in DDIM song2020DDIM scheduler. Because the rendered images $\mathbf{V}_{0}$ share common pixels in certain regions, their latents also have relevance to each other in a way indicating pixels’ move. Moreover, while being perturbed with random noise, blank spaces and flawed regions in $\mathbf{V}_{0}$ can be further filled with randomness, providing video model with more possibilities to generate and correct them.
Video Generation
After camera motion inversion, noisy latents presenting camera movements are then passed through the backward process of video diffusion model, utilizing their layout controllability to guide video generation. Leveraging prior knowledge of base video model, the generation process also bestows video with rational dynamic information. In this way, explicit camera movements are injected into video diffusion model in an appropriate and training-free fashion. Starting from noisy motion latents at timestep $t_{0}$, the generation step can be represented as:
$\hat{\mathbf{V}}_{t-1}=\sqrt{\alpha_{t-1}}\left(\frac{\mathbf{V}_{t}-\sqrt{1-\alpha_{t}}\epsilon^{(t)}_{\theta}(\mathbf{V}_{t})}{\sqrt{\alpha_{t}}}\right)+\sqrt{1-\alpha_{t-1}-\sigma^{2}_{t}}\epsilon^{(t)}_{\theta}(\mathbf{V}_{t})+\sigma_{t}\epsilon,\quad t\in\left[1,t_{0}\right].$ (5)
Here we use DDIM as sampling scheduler song2020DDIM and $\epsilon_{\theta}$ denotes the video model for noise prediction. $\sigma_{t}$ determines whether the denoising process is deterministic or probabilistic, we set $\sigma=1$ to encourage diversity of generation results.
Trade-off Between Fidelity and Diversity
Leveraging noise prior guidance in diffusion model could lead to trade-off problem between generation’s fidelity and diversity meng2021sdedit ; kim2022diffusionclip ; hou2024high , where results that hold more faithfulness towards guidance tend to decline in generation quality or diversity. In this task, similar circumstance also occurs as model is required to receive guidance from a series of imperfect renderings while generate a reasonable and dynamic video. The key factor to balance the trade-off problem lies in the choice of $t_{0}$. When larger $t_{0}$ applied, generation bears more resemblance to original guidance $\mathbf{V}_{0}$, yet lacks of rationality and dynamics to be an appealing video. Instead, smaller $t_{0}$ leads to well-generated video, but is less aligned with desired camera motion. In our experiments, we find larger $t_{0}$ works better for motions with moderate intensity, and for those with relatively drastic move, smaller $t_{0}$ shows preferable performance.
4 Experiments
4.1 Experimental Settings
Implementation Details
The major results presented in this paper are grounded on zeng2023PixelDance . To ensure a fair comparison with other state-of-the-art methods, we employ SVD as base model for quantitative evaluation. Inversion and generation steps among all methods are set to 25. Our method doesn’t require any additional training utilizing camera trajectories.
Evaluation Details
We compare CamTrol with state-of-the-art works: AnimateDiff guo2023animatediff , MotionCtrl wang2023motionctrl and CameraCtrl he2024cameractrl , all three methods are finetune on SVD using specific data. In quantitative evaluation, FVD unterthiner2018FVD , FID heusel2017FID and IS saito2020IS are used to assess video generation quality, while CLIPSIM wu2021CLIPSIM quantifies the similarity between generated video and input prompt. With regard to the accuracy of camera motion, we adopt ParticleSFMzhao2022particlesfm and produce estimated camera trajectories from generated videos, with the use of Absolute Trajectory Error(ATE) measuring their differences compared to ground truth. Relative Pose Error(RPE) is calculated to assess between consecutive frames how well the relative motions match expected ones including their transition(RPE-T) and rotation part(RPE-R). Settings of evaluation dataset follow those established in MotionCtrl wang2023motionctrl , but include more complex trajectories. Specifically, we extract 51 distinct trajectories from RealEstate10krealestate10k , each paired with 10 prompts mentioned in MotionCtrl wang2023motionctrl , resulting in 510 samples in total for assessment. As AnimateDiff guo2023animatediff lacks the ability to handle complicated trajectories, it is only included in qualitative analysis part evaluating with 8 basic camera motions. Quantitative comparisons contain AnimateDiff guo2023animatediff can be found in Appendix A.3.
4.2 Comparisons with State-of-the-art Methods
Quantitative Evaluation
Quantitative evaluations are shown in Table  1. Building upon the same base model as MotionCtrl wang2023motionctrl and CameraCtrl he2024cameractrl , our method demonstrates superior performance across all quantitative metrics concerning both video quality and camera motion accuracy. Despite being training-free, CamTrol outperforms those that rely heavily on extensive finetuning and large-scale annotated data, attaining the lowest score in ATE, RPE-T and RPE-R. Benefit from explicit camera movements modeling and motion inversion, our method is capable to produce videos align best with those tricky trajectories while maintaining their visual qualities.
Table 1: Quantitative comparisons.
Method Video Quality Motion Accuracy
FVD $\downarrow$ FID $\downarrow$ IS $\uparrow$ CLIP-SIM $\uparrow$ ATE $\downarrow$ RPE-T $\downarrow$ RPE-R $\downarrow$
MotionCtrl wang2023motionctrl 3576.40 239.10 7.58 0.2933 4.006 1.086 0.106
CameraCtrl he2024cameractrl 2922.99 243.98 8.07 0.2915 4.200 1.487 0.080
SVD+CamTrol 2832.59 227.36 8.07 0.3100 3.917 0.947 0.017
Figure 3: Qualitative comparisons with finetuned methods. CamTrol generates videos that adhere to desired camera movements, without compromising their dynamics and rationality.
Qualitative Analysis
Qualitative comparisons of different methods are illustrated in Fig. 3. All methods generate videos align well with desired camera motion, however, the constraints imposed by camera trajectories are sometimes too strong, that output videos have to compromise their quality and rationality to satisfy these movements. Some examples that demonstrate this are results of Zoom Out and Roll Anticlockwise, which looks more like flat images being zoomed or rolled on 2D plane instead of videos recorded in 3D spaces. The reason might lies on the limitations of proprietary dataset used for finetuning realestate10k ; yu2023mvimgnet , where most scenes are static, making pretrained model forget about knowledge of dynamic contents. In contrast, CamTrol makes full use of the powerful dynamics in base model and generate plausible videos with camera perspective change utilizing layout prior guidance, handles these circumstances gracefully. More comparisons at Appendix A.3.
4.3 Ablation Study
Figure 4: Comparison with base model. Controlling camera motion via prompt engineering doesn’t work at most times. Instead, CamTrol offers robust control towards video’s camera movement in a training-free manner.
Figure 5: Effectiveness of layout prior. Layout prior guidance compensates for the vacancies(Left) and static contents(Right) caused by point cloud rendering.
Comparison to Base Model
To demonstrate that changes of camera motion are attributed to our method rather than the innate capability of video model, we conduct ablation study to assess its effectiveness. We add prompts describing certain camera moves(e.g. zooms out), letting video model understand by itself. The results are shown in Fig. 4. It could be observed that even provided with prompts indicating how camera should move, base model fail to produce correct results. Instead, CamTrol  is able to implement designated motion control without any instructions from text prompts.
Effectiveness of Layout Prior
We employ ablation study to validate the effectiveness of layout prior guidance, illustrating its necessity from two aspects: the completeness of vacancies and dynamic of generated video. In Fig. 5, we showcase frames before and after noise prior guidance. With camera pose changes, there appears regions unfilled in point cloud and causes blank spaces in rendered images(left part); Besides, due to the static nature of point cloud, rendered images remain stationary(right part). Noise layout prior could compensate for these flaws, finally produce videos with inpainted vacancies and rationalized dynamics.
Effect of Timestep $t_{0}$
$t_{0}$ is a crucial factor that influences the trade-off between generated video’s diversity and its faithfulness to camera motions requirements. To investigate its effect on the output, we conduct experiments with various $t_{0}$ values, relevant results are shown in Fig. 7. As illustrated, videos generated with larger $t_{0}$ tend to conform better to camera motion requirements, but suffer from decrease in dynamics; On the contrary, smaller $t_{0}$ leads to more plausible generations but fails to meet camera’s requirements, as latents at these timesteps carry more randomness.
Generalization to Diverse Base Models
Our proposed CamTrol can be seamlessly plugged and played with most pretrained video diffusion models to achieve training-free camera control. We present visual results showcasing its applications based on different video base models, including SVD SVD and VideoFusion luo2023videofusion , in Fig. 8. Our approach remains effective applied to alternative video base models, demonstrating its strong robustness and generalization ability. More results with other base models can be found in Appendix A.5.
4.4 Further Applications
Figure 6: Effect of $t_{0}$. Smaller $t_{0}$ encourages
dynamics while larger $t_{0}$ preserves camera
movements (Pedestal Down).
Figure 7: Camera control at different scales. CamTrol supports camera movements over various scales.
Figure 8: Applied onto SVD SVD and VideoFusion luo2023videofusion . CamTrol could be plugged and played with most video diffusion models to offer camera movement control.
3D Rotation Videos
One of the most advantages of our method is it can generate videos rotating around some objects and produce outputs similar to 3D generation models sv3d ; im3d . While these 3D models need large-scale training on 3D dataset and could only handle inputs in specific styles, our approach is able to deal with any type of images and achieve this in a completely zero-shot manner. Some results are shown in 1. We offer more results on multi-view synthesis in Appendix A.2.
Hybrid and Complex Camera Movements
By combining different basic camera trajectories, CamTrol can support hybrid camera movements and endow generated video with cinematic charm. Besides this, explicit motion modeling also equips CamTrol with the abilities to support trajectories containing precise coordinates, which means it can generate videos presenting any complicated camera movement. Results about hybrid and complex motions are shown in Fig. 1 and Appendix A.2.
Camera Motion at Different Scales
CamTrol supports camera movements at controllable scales. We provide some results in Fig. 7. By specifying different magnitudes of camera’s extrinsic matrix within point cloud spaces, rendered images will exhibit varying degrees of motion, leading to videos with distinct scales of camera movements. This further demonstrate the powerful controllability of CamTrol, and provides a new pathway for video’s customized camera control.
5 Conclusion
In this paper, we propose a training-free and robust method CamTrol to offer camera control for off-the-shelf video diffusion models. It consists of two-stage procedure including explicit camera motion modeling in 3D point cloud space and video generation utilizing layout prior of noisy latents. Compared to previours work, CamTrol  does not require any additional finetuning on camera-annotated datasets, or self-supervised training via data augmentation, instead, it could be plugged and played with most video diffusion models to generate camera controllable videos with single image or text prompt as input. Comprehensive experimental results demonstrate the effectiveness and robustness of CamTrol for controlling camera motion for videos. Besides, we show that CamTrol produces impressive results in generating videos with 3D rotations, complicated trajectories, and different moving scales.
Appendix A Appendix
A.1 Definitions of Basic Camera Motions
We refer to the terminology in cinematography to describe different camera motions, definitions of each type are detailed in Table 2.
Table 2: Definitions of basic camera motions.
Camera Motion Directions Definition
Zoom In Camera moves towards or away from a subject.
Out
Tilt Up Rotating the camera vertically from a fixed position.
Down
Pan Left Rotating the camera horizontally from a fixed position.
Right
Pedestal Up Moving a camera vertically in its entirety.
Down
Truck Left Moving a camera horizontally in its entirety.
Right
Roll Clockwise Rotating a camera in its entirety in a horizontal manner.
Anticlockwise
Rotate Clockwise Moving a camera around a subject.
Anticlockwise
Hybrid Arbitrary Combination of other motions.
A.2 More Results on Camera Control
In this section, we showcase additional qualitative results of CamTrol on various camera movements. Fig. 9 shows video frames with camera move Rotate Anticlockwise and Rotate Clockwise. These outputs share sort of similarity with outputs of 3D generation models, as they all exhibit in a turning-table like way, which camera rotates around some objects. The difference here is that 3D model, as only trained on specific datasets, could only generate outputs in certain styles, e.g., single static object with no background. Instead, our model could handle arbitrary image as input, and generate a rotating-around video with proper dynamics. From this aspect, our method could be seen as a infinite source of attaining 3D data. And by utilizing our method with stronger backbones, video foundation models could truly become the largest source of 3D data as it should be.
Besides, we provide additional results on hybrid motions and basic motions, which is illustrated in Fig. 10, Fig. 11 and Fig. 13, respectively. We also present pre-loaded complex camera motions in Fig. 12.
A.3 More Comparisons to State-of-the-Art Methods
More comparison results with state-of-the-art methods are shown in Fig. 14 and Fig. 15. Compared to other works, our method matches well with input texts and camera requirements, meanwhile has wide-ranging diversities on generated contents.
We also present quantitative comparisons that include AnimateDiff guo2023animatediff as part of the evaluation. As AnimateDiff guo2023animatediff lacks the ability to handle complicated trajectories, experiments are conducted focusing on 8 basic trajectories described in A.1. To expand the diversity of the models evaluated, we employ CamTrol on zeng2023PixelDance for assessment. With regard to basic trajectories, we apply UCF-101 soomro2012UCF as the reference of calculating FID and FVD, from where we randomly sample 125 prompts per camera motion type and form 1000 samples in total for assessment. Relevant results are exhibited in Table 3.
Table 3: Quantitative comparisons on basic trajectories.
Method Video Quality
FVD $\downarrow$ FID $\downarrow$ IS $\uparrow$ CLIP-SIM $\uparrow$
AnimateDiff guo2023animatediff 1582.93 73.79 13.39 0.3196
MotionCtrl wang2023motionctrl 1768.25 78.26 14.60 0.3143
CameraCtrl he2024cameractrl 1383.68 76.18 14.04 0.3123
CamTrol+zeng2023PixelDance 1134.71 69.70 15.87 0.3253
A.4 Ablation on Depth Optimization
Ablation results relevant to the effectiveness of depth optimization are presented in Table 4. While it affects little on CLIPSIM, depth optimization exhibits benefits on other video quality assessments(FVD, FID, IS) and camera trajectory criteria(ATE, RPE). Experiments are conducted on complicated 10 trajectories extracted from RealEstate10k realestate10k , each with 10 prompts mentioned in MotionCtrl wang2023motionctrl . We assume that depth optimization offers better alignment between adjacent views rendered from point cloud, which benefits both perceptual quality and camera motion accuracy of generated videos.
Table 4: Ablation on depth coefficient optimization.
[…]
w/o optimization 5033.44 219.01 5.67 0.2949 4.093 1.131 0.046
w/ optimization 4977.88 218.36 5.72 0.2947 3.906 0.978 0.049
A.5 More Results on other Base Video Models
Our methods could be plugged and played with arbitrary video diffusion models to achieve camera control. We showcase its performance based on two alternative video generation backbones including SVD SVD and VideoFusion luo2023videofusion , relevant results are exhibited in Fig. 16.
Figure 9: Results with camera move Rotate Anticlockwise and Rotate Clockwise. Our method could generated 3D rotation-like videos with dynamics.
Figure 10: Results with camera move Hybrid I. Here we combine Zoom Out, Truck Left, Pedestal Up, Pan Right and Tilt Down together, form a cinematic like effect on generated videos.
Figure 11: Results with camera move Hybrid II. Here the hybrid motion is defined as Zoom In first, then Pedestal Up.
Figure 12: Results with complex camera motions.
Figure 13: Results with basic camera motions.
Figure 14: Comparison to state-of-the-art methods on Truck Left, Truck Right, Pedestal Up, Pedestal Down.
Figure 15: Comparison to state-of-the-art methods on Zoom In, Zoom Out, Roll Anticlockwise, Roll Clockwise.
Figure 16: Results with other video generation base models including SVD SVD and VideoFusion luo2023videofusion . Our method works well with these backbones.